# Vector-ICL: In-context Learning with Continuous Vector Representations

### v1

@@ Abstract
Large language models (LLMs) have shown remarkable in-context learning (ICL) capabilities on textual data. We explore whether these capabilities can be extended to continuous vectors from diverse domains, obtained from black-box pretrained encoders. By aligning input data with an LLM’s embedding space through lightweight projectors, we observe that LLMs can effectively process and learn from these projected vectors, which we term Vector-ICL. In particular, we find that pretraining projectors with general language modeling objectives enables Vector-ICL, while task-specific finetuning further enhances performance. In our experiments across various tasks and modalities, including text reconstruction, numerical function regression, text classification, summarization, molecule captioning, time-series classification, graph classification, and fMRI decoding, Vector-ICL often surpasses both few-shot ICL and domain-specific model or tuning. We further conduct analyses and case studies, indicating the potential of LLMs to process vector representations beyond traditional token-based paradigms. footnotetext: Code will be available at: https://github.com/EvanZhuang/vector-icl.
@@ 1 Introduction
In-context learning (ICL) has emerged as a powerful paradigm in large language models (LLMs), allowing generalization from limited examples within a given context [9, 41]. By providing demonstrations in the context during inference, ICL allows models to adapt to new tasks and formats without the need for retraining. However, since LLMs are trained on discrete natural language tokens, ICL is generally learned and used through natural language, limiting its applicability to non-textual data.
We explore whether LLMs can perform ICL directly on continuous vectors, a capability that could dramatically expand their applicability. Many data modalities, such as sensor readings, financial time series, or scientific measurements, lack a natural text representation. Moreover, even for text data, information like numbers might be better represented via continuous vectors than token sequences.
@@
Our investigation begins with the task of text reconstruction, where we assess whether LLMs can recover information encoded in text embedding. This serves as a proof-of-concept for Vector-ICL, showing that LLMs can indeed extract meaningful information from projected continuous vectors. We then turn to the more complex challenge of arithmetics. Although state-of-the-art LLMs can solve Olympiad mathematical problems [51, 41], they struggle with precise number processing due to the limitations inherent in their tokenization schemes. Our results demonstrate that Vector-ICL offers a more effective approach for function approximation, particularly for large numbers that span multiple tokens, potentially opening new avenues for enhancing LLMs’ numerical reasoning capabilities.
@@ 2 Related work
ICL has empirically shown strong performance in diverse natural-language processing tasks with very few demonstrations [9, 41]. In modern LLMs with long context windows, ICL has even shown performance improvements as the number of demonstrations grows to hundreds or even thousands, sometimes outperforming finetuning [1, 25, 7]. Empirically, different factors play key roles in ICL. In smaller LLMs, ground-truth demonstrations are not required for in-context learning, while other factors such as the label space, input text distribution, and overall sequence format play an important role [35]. Moreover, these LLMs can sometimes achieve strong performance even when demonstrations are intentionally irrelevant or even pathologically misleading [53]. More recently, Wei et al. [54] characterize these behaviors of LLMs with respect to model size, and show that larger language models perform in-context learning differently in the presence of flipped or semantically unrelated labels. Orthogonally, different works find ways to improve ICL, e.g. by including explanations [22], or chaining ICL calls [36]. ICL has shown some success in multimodal models [57, 21] or when applied to tabular data [64].
@@
Many works have investigated ICL and found that it is able to learn linear models [3, 62], discrete functions [8], and more general algorithms [27]. Some works have explicitly connected ICL in specific settings to implementing optimization steps analogous to gradient descent [30, 52, 2] and higher-order optimization methods [11, 15, 16, 62]. A complementary direction aims to establish statistical complexity and generalization bounds of in-context learning in transformers [5, 27, 55, 56]. Finally, one recent work suggests that ICL may arise from parallel structures in pretraining data [10].
@@
In contrast to the emergent ICL capabilities of LLMs, existing works have also studied how to explicitly improve ICL.  Min et al. [34] propose MetaICL, a meta-training framework for finetuning pretrained LLMs to perform in-context learning on a large and diverse collection of tasks. In the tabular domain, TNP [39] and PFNs [37] train transformer models to perform in-context prediction for a family of functions, which allows in-context generalization to unseen functions after training. Zhao et al. [63] also propose meta-learning transformers to in-context learn group preferences, serving as an in-context learned reward model that adapts to diverse group preferences.
@@ 3 Method: vector context via embedding projection
Figure 3: Main results: LLMs can perform Vector-ICL ($\uparrow$ = better). We show that training the embedding projector with a simple next-token prediction objective enables Vector-ICL. Fine-tuning the projector on downstream tasks further improves the use of continuous context, surpassing the performance of few-shot ICL and domain-specific baseline models. The study begins with text reconstruction to assess LLMs’ ability to interpret box token embeddings, followed by function regression to evaluate reasoning capabilities. We then demonstrate Vector-ICL’s effectiveness and applicability across various downstream tasks, including text classification, summarization, time-series classification, graph classification, and brain fMRI decoding & classification. Results in each panel are averaged over different encoders and LLMs for the diverse tasks we study; error bars show 95% confidence intervals.
@@ 3.1 Embedding projection
To perform Vector-ICL, an input must be transformed into a vector context through an embedding projection. Given a dataset $\mathcal{X}=\{x_{i}\}_{i=1}^{n}$, we assume the existence of an encoder $f_{\textrm{enc}}$, that transforms the data into an abstract representation (alternatively, the raw data may already be a continuous vector). The encoded embeddings, $f_{\textrm{enc}}(x)$, are then projected into box tokens, denoted as $\Box_{x}$. Throughout the paper, we will use the terms “box tokens” and “projected embeddings” interchangeably. For the decoding step, we employ an LLM, referred to as $f_{\textrm{llm}}$, to generate the output given the prompts.
@@
The MLP follows the architecture of the MLP block in Llama [50], with an additional input projection layer to map the input dimension $\mathbb{R}^{d_{\mathrm{enc}}}$ to output dimension $\mathbb{R}^{d_{\mathrm{dec}}}$.
@@ 3.3 Training the embedding projectors
The finetuning process is shown in Fig. 2(b). It utilizes additional structured prompts and trains with task-specific datasets. Similarly, the input is first mapped into the embedding space and projected into $\Box$ tokens. They are then inserted into structured prompts, while the LLM is trained with conditional generation loss given those prompts.
@@ 4 Experimental setup
To pretrain our text projectors, we leverage the WikiText-103 [33] dataset, consisting of over 100 million tokens from verified high-quality Wikipedia articles. This smaller language modeling corpus is chosen for its suitability to the lightweight nature of our projectors. The pretraining process is illustrated in Fig. 2(a), where text snippets are divided at random sentence-end points. The first half is embedded and projected, while a next-token generation loss is applied to the second half.
@@
We investigate LLMs’ ability to decode original text from projected embeddings using two datasets: Parallel Sentence Talks [49] and Quora [48]. These datasets consist of concise text pieces that convey clear meaning. Projectors are trained on the training sets of both datasets, and performance is evaluated using the BLEU score [43, 44].
@@
Following the classification tasks, we explore Vector-ICL’s capability in summarizing text based on the projected embeddings. The datasets, encoders, LLMs, and prompt templates can be found in Table 1. Performance is evaluated using RougeL [28].
@@
We also extend our approach to the unconventional task of molecule captioning, using molecule sequence-caption pairs from the Language + Molecules-24 (LPM24) [14] dataset. A sample molecule-caption pair looks like the following:
@@
We take the output of the last time step from Chronos-base [4] as the time-series’ representation. We use the base encoder with trained classification head as the baseline and we measure the prediction performance with accuracy.
@@
We use Graphormer [59] as the encoder model, specifically the one that was pretrained on quantum chemistry graph datasets [19]. Since the down-stream, ogbg-molhiv [18], is a molecule property prediction dataset, and with strong class imbalance (3-4% positive classes), we finetune the encoder on the training set to provide meaningful baselines and embeddings. We take the output prior to the classification layer of the Graphormer as the graph embedding. Weighted sampling is adopted in the finetuning of both the baseline Graphormer and the embedding projector to yield meaningful predictions. We use the finetuned Graphormer as our baseline and use the F1 score as the performance metric.
@@
Table 2: We compare Vector-ICL with few-shot ICL and soft prompt tuning on the five sentiment analysis datasets. Details can be found in Sec. A.2.
Method Rotten Tomatoes SST2 IMDB Emotion Financial Phrasebank
V-ICL (pretrained) 80.60 78.90 95.04 41.20 60.72
Few-shot ICL 87.31 91.74 93.50 55.20 71.78
Soft prompt [26] 93.24 96.21 95.26 74.15 78.22
V-ICL (finetuned) 88.80 98.16 97.28 85.20 81.68
@@ 5 Results: unlocking versatile applications across modalities
Table 3: We compare the performance of Vector-ICL, few-shot ICL, and soft prompt tuning across two text summarization datasets. Details can be found in Sec. A.2.
Method XSum XLSum
V-ICL (pretrained) 15.25 15.89
Few-shot ICL 19.53 19.41
Soft prompt 12.84 12.70
V-ICL (finetuned) 20.08 20.49
@@ 6.1 The Impact of Encoder Quality on Vector-ICL Performance
Our analysis focuses on text classification as the downstream task. We examine the correlation between encoder rankings on the reconstruction task and their corresponding rankings on the classification task. This correlation test is performed across 5 datasets and 3 LLMs, resulting in 15 configurations.
@@ 6.2 Case study: what has been learned in the projections?
Fig. 4(b) provides a visualization of the normalized Euclidean distances between projected embeddings, where the x and y axes correspond to 1024 different numbers uniformly sampled from 00 to $101e$ 10 and the color intensity reflects the distance between each pair. Several key patterns emerge from this visualization that offer insights into what the projector has learned.
@@
Another notable feature of the distance matrix is the block structure that emerges. This pattern comes from the way we construct the numerical embeddings and it is likely beneficial for the LLM to understand numerical inputs.
Figure 4: Key Insights from Encoders and Projections. (a) The impact of encoders’ information-preserving capabilities on downstream task performance with Vector-ICL. We observe that there is a positive correlation, suggesting that an encoder’s ability to retain information may be a strong predictor of its effectiveness in Vector-ICL. (b) The Euclidean distance matrix between uniformly sampled numbers after projection, using the projector trained with Qwen2-7B as an example. The projected number embeddings exhibit block-diagonal patterns, indicating structures and sensible distance measures. (c) Analyzing LLM’s understanding of projected brain fMRI embeddings after only unsupervised pretraining. We categorize the underlying questions related to the text and measure the mean accuracy for each category, highlighting the LLM’s ability to interpret the embeddings, with only the next token prediction pretraining.
6.3 Case study: what information was perceived from projected brain fmri?
Fig. 4(c) illustrates the mean accuracy achieved in decoding different categories of brain activity based on fMRI data. The categories, ranging from ”Physical Actions and Movements” to ”Conflict, Urgency, and Change,” represent diverse cognitive and perceptual domains. The grouping and corresponding questions are listed in Table 5.
@@ 7 Discussion
In this study, we explored a variety of settings: utilizing different encoders, LLM architectures, modalities, and datasets. Our results demonstrate that LLMs are capable of performing Vector-ICL on both language and non-language inputs. However, our experiments did not cover all possible combinations of these variables. There are still many unexplored areas, such as additional modalities, tasks, and encoder-decoder configurations, that could further benefit from Vector-ICL. Also, we only experimented with single-token encoders, while there exist encoders that produce variable-sized embeddings, that can potentially be more powerful and flexible. We leave this extensive exploration for future research to fully understand the broader applicability and limitations of our approach across diverse domains.

### v2

@@ Abstract
Large language models (LLMs) have shown remarkable in-context learning (ICL) capabilities on textual data. We explore whether these capabilities can be extended to continuous vectors from diverse domains, obtained from black-box pretrained encoders. By aligning input data with an LLM’s embedding space through lightweight projectors, we observe that LLMs can effectively process and learn from these projected vectors, which we term Vector-ICL. In particular, we find that pretraining projectors with general language modeling objectives enables Vector-ICL, while task-specific finetuning further enhances performance. In our experiments across various tasks and modalities, including text reconstruction, numerical function regression, text classification, summarization, molecule captioning, time-series classification, graph classification, and fMRI decoding, Vector-ICL often surpasses both few-shot ICL and domain-specific model or tuning. We further conduct analyses and case studies, indicating the potential of LLMs to process vector representations beyond traditional token-based paradigms. footnotetext: Code is available at: https://github.com/EvanZhuang/vector-icl.
@@ 1 Introduction
In-context learning (ICL) has emerged as a powerful paradigm in large language models (LLMs), allowing generalization from limited examples within a given context (Brown et al., 2020; OpenAI, 2023). By providing demonstrations in the context during inference, ICL allows models to adapt to new tasks and formats without the need for retraining. However, since LLMs are trained on discrete natural language tokens, ICL is generally learned and used through natural language, limiting its applicability to non-textual data.
We explore whether LLMs can perform ICL directly on continuous vectors, a capability that could dramatically expand their applicability. Many data modalities, such as sensor readings, financial time series, or scientific measurements, lack a natural text representation. Moreover, even for text data, information like numbers might be better represented via continuous vectors than tokens.
@@
Our investigation begins with the task of text reconstruction, where we assess whether LLMs can recover information encoded in text embedding. This serves as a proof-of-concept for Vector-ICL, showing that LLMs can indeed extract meaningful information from projected continuous vectors. We then turn to the more complex challenge of arithmetics. Although state-of-the-art LLMs can solve Olympiad mathematical problems (Trinh et al., 2024; OpenAI, 2023), they struggle with precise number processing due to the limitations inherent in their tokenization schemes. Our results demonstrate that Vector-ICL offers a more effective approach for function approximation, particularly for large numbers that span multiple tokens, potentially opening new avenues for enhancing LLMs’ numerical reasoning capabilities.
@@ 2 Related work
ICL has empirically shown strong performance in diverse natural-language processing tasks with very few demonstrations (Brown et al., 2020; OpenAI, 2023). In modern LLMs with long context windows, ICL has even shown performance improvements as the number of demonstrations grows to hundreds or even thousands, sometimes outperforming finetuning (Agarwal et al., 2024; Li et al., 2023a; Bertsch et al., 2024). Empirically, different factors play key roles in ICL. In smaller LLMs, ground-truth demonstrations are not required for in-context learning, while other factors such as the label space, input text distribution, and overall sequence format play an important role (Min et al., 2022b). Moreover, these LLMs can sometimes achieve strong performance even when demonstrations are intentionally irrelevant or even pathologically misleading (Webson & Pavlick, 2022). Flan-T5 (Chung et al., 2024) revealed that instruction tuning improves few-shot learning by helping LLMs better utilize in-context examples in an encoder-decoder architecture.More recently, Wei et al. (2023) characterize these behaviors of LLMs with respect to model size, and show that larger language models perform in-context learning differently in the presence of flipped or semantically unrelated labels. Orthogonally, different works find ways to improve ICL, e.g. by including explanations (Lampinen et al., 2022), or chaining ICL calls (Morris et al., 2023). ICL has shown some success in multimodal models (Wu et al., 2024; Jiang et al., 2024) or when applied to tabular data (Zhao et al., 2024).
@@
Many works have investigated ICL and found that it is able to learn linear models (Akyürek et al., 2022; Zhang et al., 2023), discrete functions (Bhattamishra et al., 2023), and more general algorithms (Li et al., 2023b). Some works have explicitly connected ICL in specific settings to implementing optimization steps analogous to gradient descent (Mahankali et al., 2023; Von Oswald et al., 2023; Ahn et al., 2024) and higher-order optimization methods (Dai et al., 2023; Fu et al., 2023; Giannou et al., 2024; Zhang et al., 2023). A complementary direction aims to establish statistical complexity and generalization bounds of in-context learning in transformers (Bai et al., 2024; Li et al., 2023b; Wies et al., 2024; Wu et al., 2023). Finally, one recent work suggests that ICL may arise from parallel structures in pretraining data (Chen et al., 2024).
@@
In contrast to the emergent ICL capabilities of LLMs, existing works have also studied how to explicitly improve ICL.  Min et al. (2022a) propose MetaICL, a meta-training framework for finetuning pretrained LLMs to perform in-context learning on a large and diverse collection of tasks. In the tabular domain, TNP (Nguyen & Grover, 2022) and PFNs (Müller et al., 2021) train transformer models to perform in-context prediction for a family of functions, which allows in-context generalization to unseen functions after training. Zhao et al. (2023) also propose meta-learning transformers to in-context learn group preferences, serving as an in-context learned reward model that adapts to diverse group preferences.
@@ 3 Method: vector context via embedding projection
Figure 3: Main results: LLMs can perform Vector-ICL ($\uparrow$ = better). We show that training the embedding projector with a simple next-token prediction objective enables Vector-ICL. Even with only unsupervised pretraining, Vector-ICL matches or outperforms traditional few-shot ICL on 4 out of 6 tasks where direct comparison is possible. Fine-tuning the projector on downstream tasks further enhances the use of continuous context, consistently surpassing both few-shot ICL and specialized task-tuned baselines (soft-prompt for text, tuned encoders for non-text).The study begins with text reconstruction to assess LLMs’ ability to interpret box token embeddings, followed by function regression to evaluate reasoning capabilities. We then demonstrate Vector-ICL’s effectiveness and applicability across various downstream tasks, including text classification, summarization, time-series classification, graph classification, and brain fMRI decoding & classification. Results in each panel are averaged over different encoders and LLMs for the diverse tasks we study; error bars show 95% confidence intervals.
@@ 3.1 Embedding projection
Vector-ICL requires transforming inputs into vector contexts through an embedding projection. Given a dataset $\mathcal{X}=\{x_{i}\}_{i=1}^{n}$, we assume the existence of an encoder $f_{\textrm{enc}}$, that transforms the data into an abstract representation (alternatively, the raw data may already be a continuous vector). The encoded embeddings, $f_{\textrm{enc}}(x)$, are then projected into box tokens, denoted as $\Box_{x}$. Throughout the paper, we will use the terms “box tokens” and “projected embeddings” interchangeably. For decoding, we use a language model $f_{\textrm{llm}}$ that generates outputs based on the provided prompts.
@@
The MLP follows the architecture of the MLP block in Llama (Touvron et al., 2023), with an additional input projection layer to map the input dimension $\mathbb{R}^{d_{\mathrm{enc}}}$ to output dimension $\mathbb{R}^{d_{\mathrm{dec}}}$.
@@ 3.3 Training the embedding projectors
The finetuning process is shown in Fig. 2(b). It utilizes additional structured prompts and trains with task-specific datasets. Similarly, the input is first mapped into the embedding space and projected into $\Box$ tokens. They are then inserted into structured prompts, while the projector is trained with conditional generation loss given those prompts.
@@ 4 Experimental setup
Baselines
We evaluate Vector-ICL in two distinct settings: with and without task-specific tuning. For textual tasks, we compare against few-shot ICL and soft prompt tuning (Li & Liang, 2021). We choose soft prompt tuning as our primary baseline because it represents a similarly lightweight adaptation approach - both methods introduce a small number of trainable parameters while keeping the base LLM frozen. Like our projector, soft prompts modify how the LLM processes inputs without changing its internal weights. This makes it a fair comparison point for assessing whether Vector-ICL’s benefits come from the continuous vector representations themselves rather than just the additional training. For non-textual tasks, where soft prompts cannot be directly applied, we compare against cross-modal few-shot ICL and tuned encoders. In the non-textual domain, ICL inputs are represented either as numeric sequences (for time-series and brain fMRI data) or as textual descriptions (for graph edge lists and node features).
@@
To pretrain our text projectors, we leverage the WikiText-103 (Merity et al., 2016) dataset, consisting of over 100 million tokens from verified high-quality Wikipedia articles. This smaller language modeling corpus is chosen for its suitability to the lightweight nature of our projectors. The pretraining process is illustrated in Fig. 2(a), where text snippets are divided at random sentence-end points. The first half is embedded and projected, while a next-token generation loss is applied to the second half.
@@
We investigate LLMs’ ability to decode original text from projected embeddings using two datasets: Parallel Sentence Talks (Tiedemann, 2012) and Quora (Thakur et al., 2021). These datasets consist of concise text pieces that convey clear meaning. Projectors are trained on the training sets of both datasets, and performance is evaluated using the BLEU score (Papineni et al., 2002; Post, 2018).
@@
Following the classification tasks, we explore Vector-ICL’s capability in summarizing text based on the projected embeddings. The datasets, encoders, LLMs, and prompt templates can be found in Table 1. Performance is evaluated using RougeL (Lin, 2004).
@@
We also extend our approach to the unconventional task of molecule captioning, using molecule sequence-caption pairs from the Language + Molecules-24 (LPM24) (Edwards et al., 2024) dataset. A sample molecule-caption pair looks like the following:
@@
We take the output of the last time step from Chronos-base (Ansari et al., 2024) as the time-series’ representation. We use the base encoder with trained classification head as the baseline and we measure the prediction performance with accuracy.
@@
We use Graphormer (Ying et al., 2021) as the encoder model, specifically the one that was pretrained on quantum chemistry graph datasets (Hu et al., 2021). Since the down-stream, ogbg-molhiv (Hu et al., 2020), is a molecule property prediction dataset, and with strong class imbalance (3-4% positive classes), we finetune the encoder on the training set to provide meaningful baselines and embeddings. We take the output prior to the classification layer of the Graphormer as the graph embedding. Weighted sampling is adopted in the finetuning of both the baseline Graphormer and the embedding projector to yield meaningful predictions. We use the finetuned Graphormer as our baseline and use the F1 score as the performance metric.
@@
Table 2: Comparison of Vector-ICL, few-shot ICL, and soft prompt tuning across various sentiment analysis and summarization datasets. Details can be found in Sec. A.3.
Method Sentiment Analysis Summarization
Rotten Tomatoes SST2 IMDB Emotion Financial Phrasebank XSum XLSum
V-ICL (pretrained) 80.60 78.90 95.04 41.20 60.72 15.25 15.89
Few-shot ICL 87.31 91.74 93.50 55.20 71.78 19.53 19.41
Soft Prompt 93.24 96.21 95.26 74.15 78.22 12.84 12.70
V-ICL (finetuned) 88.80 98.16 97.28 85.20 81.68 20.08 20.49
@@ 6.1 The Impact of Encoder Quality on Vector-ICL Performance
Our analysis focuses on text classification as the downstream task. We examine the correlation between encoder rankings on the reconstruction task and their corresponding rankings on the classification task. This analysis is performed across 5 datasets and 3 LLMs, resulting in 15 configurations.
@@ 6.2 Case study: what has been learned in the projections?
Fig. 4(b) provides a visualization of the normalized Euclidean distances between projected embeddings. Several key patterns emerge from this that offer insights into what the projector has learned.
@@
Another notable feature of the distance matrix is the block structure that emerges. The block structure reflects how numbers share similar digit patterns across the decimal places, from local to global blocks. It likely helps LLMs process numerical relationships more effectively, as it preserves the hierarchical nature of place-value notation.
Figure 4: Key Insights from Encoders, Projections, and Synthetic Data Curation. (a) Correlation between encoders’ text reconstruction performance and their downstream task effectiveness with Vector-ICL, suggesting information preservation ability predicts Vector-ICL performance. (b) Euclidean distance matrix of 1024 projected number embeddings (0 to 1e10) shows structured block-diagonal patterns, indicating meaningful numerical relationships are preserved. (c) Results from pretraining on our synthetic time-series QA dataset, which captures statistical properties through trend analysis, anomaly detection, and stability assessment. The curated QA pairs enable effective cross-modal pertaining.
6.3 Case study: can synthetic dataset be useful in cross-modal pretraining?
The cross-modal pretraining requires data-text dual pairs, which is not available in many cases. We explore how to create a synthetic dataset for such pretraining using time-series as an example. We designed a data curation pipeline that extracts meaningful statistical properties from time-series and converts them into natural language descriptions.
Our pipeline analyzes multiple statistical aspects of the time-series data. We perform trend analysis through linear regression to capture overall directional movements, anomaly detection using z-score analysis to identify unusual patterns, and temporal stability assessment via variability thresholds to characterize consistency.
For each identified property, we generate both descriptive statements and binary questions. For instance, given a time-series segment, we generate descriptive text like ”The time-series shows an upward trend” or binary questions such as ”Does this time-series contain any anomalies?”
Our experiments show that projectors pretrained on this synthetic corpus can effectively generalize to new tasks, demonstrating the feasibility of creating synthetic data for cross-modal pretraining when natural data pairs are unavailable.
@@ 7 Discussion
In this study, we explored a variety of settings: utilizing different encoders, LLM architectures, modalities, and datasets. Our results demonstrate that LLMs are capable of performing Vector-ICL on both language and non-language inputs. However, our experiments did not cover all possible combinations of these variables. There are still many unexplored areas, such as additional modalities, tasks, and encoder-decoder configurations, that could further benefit from Vector-ICL. Also, we only experimented with single-token encoders, while there exist encoders that produce variable-sized embeddings, that can potentially be more powerful and flexible. Analyzing how instruction tuning might affect the model’s ability to understand vector context would be beneficial as well. We leave this extensive exploration for future research to fully understand the broader applicability and limitations of our approach across diverse domains.